I am having a hard time with this problem: 4 1/8 - 1 3/4 If you could help, that would be appreciated. Thank you!
Answer by ivy12003(22)

First, you want to change the mixed fractions into improper fractions by multiplying the whole numbers by the denominator and adding the numerator.

4 1/8 - 1 3/4 = (4 * 8 + 1)/8 - (1 * 4 + 3)/4 = 33/8 - 7/4

Then, you want to convert the fractions so that they have a common denominator (8).

33/8 - 7/4 = 33/8 - 14/8

Next, you subtract the numerators and keep the denominators.

33/8 - 14/8 = 19/8

Finally, you change the answer into a mixed fraction.

19/8 = 2 3/8

Therefore, your complete answer would be:

4 1/8 - 1 3/4 = 33/8 - 7/4

= 33/8 - 14/8

=19/8 = 2 3/8